OpenStudy (anonymous):

f(x)=1/2x +2...how do I place this on a graph

OpenStudy (anonymous):

It is of the form y=mx+b, where m=1/2 is the slope and b=2 is the y intercept. Draw your axes. Place a mark at 2 on the y axis. From that mark, go up 1 over to the right 2 and place another mark. Draw a long straight line through those two points.

OpenStudy (anonymous):

for to find y-intercept take x=0; we get (0,y) and to find x-intercept take y=0; we get (x,0) make a stright line through those 2 points coordinats

OpenStudy (anonymous):

still trying to figure this one out

OpenStudy (anonymous):

got this so far... f(-4)= (-4)+2

OpenStudy (anonymous):

my bad f(-4)=1/2(-4)+2=

OpenStudy (anonymous):

stuck

OpenStudy (anonymous):

@tanjung presented a valid, but completely different way to do this. If you are following his way, then place 0 in for x. I don't know where the -4 is coming from.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

s0

OpenStudy (anonymous):

f(0)=1/2(0)+2=2... so y=2 when x is zero. Hence (0,2) is the point to plot.

OpenStudy (anonymous):

f(0)=1/2(0)+2=

OpenStudy (anonymous):

yes... that equals 2.

OpenStudy (anonymous):

(0,2) on x axis

OpenStudy (anonymous):

and Y??

OpenStudy (anonymous):

(x, y) is a standard ordered pair... so (0, 2) means that the x value is 0 and the y value is 2 for this one point. That would wind up on the y axis.

OpenStudy (anonymous):

my bad I meant second plotted point

OpenStudy (anonymous):

The second point comes from setting y to zero. f(x) in your case. 0=1/2(x)+2

OpenStudy (anonymous):

Subtract 2 from both sides -2=1/2(x) Multiply both sides by 2 -4=x So x=-4 when y equals zero. That gives you a point (-4, 0), which is 4 to the left on the x axis.

OpenStudy (anonymous):

got it...

OpenStudy (anonymous):

Awesome!

OpenStudy (anonymous):

It should look a little like that.

OpenStudy (anonymous):

Thank you...Had a long day and my brain is on pause...lol

OpenStudy (anonymous):

You're very welcome.